Which of the following is true about the relationship between the slopes of the lines whose equations are 8x - 9y = 5 and 9x - 8y = 1?
They are equal.
They are opposites.
They are reciprocals.

Answer by Theo(13342)

convert the equations to slope intercept form.

for 8x - 9y = 5, .....
add 9y to both sides of the equation and subtract 5 from both sides of the equation to get:
8x - 5 = 9y
switch sides to get:
9y = 8x - 5
divide both sides by 9 to get:
y = 8/9 * x - 5/9

for 9x - 8y = 1, .....
add 8y to both sides of the equation and subtract 1 from both sides of the equation to get:
9x - 1 = 8y
switch sides to get:
8y = 9x - 1
divide both sides by 8 to get:
y = 9/8 * x - 1/8

the slopes are 8/9 and 9/8
looks like they are reciprocals of each other.